OpenStudy (anonymous):

If F(x) = x + 6 and G(x) = x^4, what is G(F(x))? A. x^4(x + 6) B. (x + 6)^4 C. x^4 + 6 D. x^4 + x + 6

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[G(F(x))=G(x+6)\]and since \[G(\spadesuit)=\spadesuit^4\] you have \[G(x+6)=(x+6)^4\]